# Large Language Model Can Continue Evolving From Mistakes

### v3

@@ Appendix A Process of Collecting Corpora from Multiple Sources
Figure 4: The process of collecting corpora from Bing.
Figure 5: The process of collecting corpora from Wikipedia.
Appendix B Experimental Results of Various LLMs on QA tasks
@@
Appendix C Experimental Results with Various supplemental Training Dataset Construction Strategies
Table 6 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after being trained with a supplemental training set constructed using various strategies.
@@ Appendix C Experimental Results with Various supplemental Training Dataset Construction Strategies
Appendix D Data Division
Table 7 shows the the data division of Xiezhi and CMMLU. The multiple-choice questions are divided into three parts. The Pre-Finetuning part is utilized to train the model to answer questions in a standard format before supplemental training. The Testing part is used to construct the test set, and the remaining part is used to construct the Normative Corpus and Extractive Corpus for supplemental training.
@@ Appendix D Data Division
Table 7: The data division of Xiezhi and CMMLU.

### v4

@@ Appendix B Process of Collecting Corpora from Multiple Sources
Figure 6: The process of collecting supplementary corpus from Bing.
Figure 7: The process of collecting supplementary corpus from Wikipedia.
Appendix C Experimental Results of LLMs on QA tasks
@@
Appendix D Experimental Results with Various supplemental Training Dataset Construction Strategies
Table 5 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after being trained with supplemental training sets constructed using the Plain, Extractive, and Review strategies, as well as various data sources.
@@ Appendix D Experimental Results with Various supplemental Training Dataset Construction Strategies
Appendix E Data Division
Table 6 and Table 7 shows the the data division of Xiezhi and CMMLU. The multiple-choice questions are divided into three parts. The Pre-Finetuning part is utilized to train the model to answer questions in a standard format before supplemental training. The Testing part is used to construct the test set, and the remaining part is used to construct the Normative Corpus and Extractive Corpus for supplemental training.
@@ Appendix E Data Division
Table 6: The data division and usage of Xiezhi and CMMLU.
Dataset Total Pre-Finetuning Testing Random Replay
HotpotQA 7,000 2,000 5,000 500
GSM8K 7,000 2,000 5,000 500
Table 7: The data division and usage of HotpotQA and GSM8K, where the data used for Random Replay is included in pre-finetuning.
Appendix F Experimental Results of Multiple Rounds of CEM
We employ the identical training data construction strategy as in the first round to conduct a second round of CEM on the LLM. Table 8 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after two rounds of CEM.
Metrics Wiki Bing
CEM-N CEM-E CEM-R CEM-N CEM-E CEM-R
Acc INIT 42.93 42.93 42.93 42.93 42.93 42.93
Round1 46.56 52.02 55.18 45.91 48.46 53.86
Round2 51.69 55.20 59.93 46.37 48.74 58.33
W2R vs. INIT 24.46 26.03 28.29 22.07 23.32 27.01
vs. Round1 20.00 17.56 16.43 18.49 18.33 16.61
R2W vs. INIT 15.70 13.76 11.29 18.63 17.51 11.61
vs. Round1 14.87 14.39 11.68 18.03 18.06 12.14
Table 8: The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi task, where "v.s. INIT" indicates W2R and R2W calculated based on the initial evaluation results, and "v.s. Round1" indicates W2R and R2W calculated based on the evaluation results after one round of CEM.
Appendix G Experimental Results of Catastrophic Forgetting Analysis
Table 9 shows the complete experimental results of CuteGPT-13B on the Xiezhi, GSM8K, and HotpotQA tasks after one, two, and three rounds of CEM, respectively.
Round Strategy Task Accuracy Relative Rate
Xiezhi GSM8K HotpotQA ER AFR
INIT 40.84 13.20 80.60 1.00 1.00
Round 1 CEM-P 42.02 6.64 51.60 1.03 0.730
CEM-E 42.76 8.36 55.52 1.05 0.681
CEM-R 51.49 7.54 55.62 1.26 0.673
Round 2 CEM-P 48.68 9.94 57.84 1.19 0.723
CEM-E 49.47 9.76 60.24 1.21 0.746
CEM-R 54.23 6.88 56.68 1.33 0.703
Round 3 CEM-P 50.58 10.02 60.96 1.24 0.757
CEM-E 52.69 9.46 62.98 1.29 0.772
CEM-R 58.91 7.48 59.86 1.44 0.678
Table 9: The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi, GSM8K, and HotpotQA tasks.